Binary Fission in bacteria if very fast, sometimes doubling every 20 minutes. If bacteria can divide every 20 minutes how many bacteria would there be after 4 hours if there were originally 100?
1
Expert's answer
2020-01-24T06:09:03-0500

As bacteria divide every 20 minutes for 4 hours, the number of their division equals:

4h / 20min = 240min / 20min = 12 divisions

The number of bacteria after 12 divisions is:

N = N0 × 212 , where N0 - the initial number of bacteria.

N = 100 × 212 = 100 × 4,096 = 4,096,100 bacteria.


Answer: 4,096,100 bacteria.